# EdgeAlpha: Bringing Process Discovery to the Data Sources

### v1

@@ Abstract
Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT). The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs. Further, privacy considerations often prevent data collection at a central location in the first place. To address this challenge, this paper introduces EdgeAlpha, a distributed algorithm for process discovery operating directly on sensor nodes and edge devices on a stream of real-time event data. Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded. From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model. EdgeAlpha enables (a) scalable mining, as a node, for each event, only interacts with its predecessors and, when queried, only exchanges aggregates, i.e., partial footprint matrices, with the central location and (b) privacy preserving process mining, as nodes only store their own as well as predecessor and successor events. On the Sepsis Cases event log, for example, a node queries on average 18.7% of all nodes. For the Hospital Log, we can even reduce the overall querying to 3.87% of the nodes.
Keywords:
@@
Distributed Process Mining Event Data Stream In-Network Processing Scalability
1 Introduction
@@
In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from IoT sensors, for example, in a smart manufacturing or smart health setting [6, 13, 17]. However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [8, 15, 24]. This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [3]. Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place [14, 19, 21]. We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
@@
Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes events directly at the source (right), and only exchanges partial footprint matrices.
@@
In this paper, we introduce EdgeAlpha, a distributed algorithm for process mining based on the Alpha Miner [4]. Utilizing the constrained compute and storage resources of sensor nodes and edge devices we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location. For each event, that a node in EdgeAlpha senses, it locally stores the event as well as the predecessor and successor of this event and updates a local footprint matrix (FM) accordingly, see Figure 1. Thus, over time, each node constructs a partial FM that represents the partial view of this node on the process. To derive the complete process model, a central entity queries the nodes for their partial FM and combines these to a complete matrix.
To determine the predecessor of an event, the node that records this particular event, has to potentially query all other nodes, which can lead to large communication overheads. In this paper, we show that events of real-world processes statistically only have a few predecessors, which further only rarely change throughout the lifetime of a process. As a result, for each event, a node can very efficiently determine the predecessor, which in turn reduces the need for messaging within the network and thereby increases scalability. Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner. EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [22].
@@
Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [9] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. This observation enables EdgeAlpha to reduce the communication overhead for determining predecessor events by up to 96% compared to querying all nodes in the network. Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events. By batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes. In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor as well as a baseline that always queries the correct node immediately.
@@
  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach when querying after each event.
  • We ensure that nodes only share aggregated event data with the central entity.
The remainder of the paper is structured as follows. Section 2 introduces the necessary background. Next, Section 3 presents related work on distributed process discovery. Afterwards, we introduce the design of EdgeAlpha in Section 4. In Section 5, we evaluate EdgeAlpha. We conclude the paper in Section 6.
2 Preliminaries and Background
In this section, we provide the required preliminaries and background. We begin with definitions of direct successions and footprint matrices, and then recapitulate the Alpha Miner Algorithm [4].
@@ 2 Preliminaries and Background
Activity $b$ directly succeeds $a$ in event log $L$, i.e. $a>_{L}b$, if and only if there exists a trace $\sigma=(t_{1},\dots,t_{n})\in L$ and $j\in\{1,\dots,n-1\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
@@
Next, we recapitulate the fundamentals of the Alpha Miner before later in the paper introducing a distributed variant. First, the algorithm, see Algorithm 1, identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$. Each activity of this set corresponds to a transition in the final output model computed by the algorithm. Next, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and the end activities of these traces, respectively. Subsequently, the algorithm generates a set $X_{L}$, which comprises pairs of activity sets where each element of the first set is directly followed by every element of the second set, although not vice versa. Moreover, activities within each set do not directly succeed one another. The Alpha Miner minimizes the set of pairs, followed by the derivation of Petri net places from the minimized pairs in the next step. Finally, the algorithm adds arcs between these places and transitions, forming the resulting Petri net.
Algorithm 1 Alpha Miner
1:$T_{L}~{}=\{t\in T~{}|~{}\exists~{}\sigma\in L:~{}~{}t\in\sigma\}$
2:$T_{I}~{}=\{first(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
3:$T_{O}=\{last(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
[…]
6:$P_{L}=\{p_{(A,B)}~{}|~{}(A,B)\in Y_{L}\}~{}\cup~{}\{i_{L},o_{L}\}$
[…]
8:$\alpha(L)=(P_{L},~{}T_{L},~{}F_{L})$
As part of the calculation of $X_{L}$ in the Alpha Miner, it first computes a so-called footprint matrix (FM). Essentially, the FM serves to store information regarding the direct succession relationships between activities within the event log. For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is element of $\{0,\dots,|T_{L}|-1\}$. This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
@@
For our design, we additionally define partial footprint matrices that contain only one specific column. For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
3 Related Work
Van der Aalst [2] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log. Here, models rather than FMs are created in a distributed manner. Evermann and Assadipour [11] leverage Map-Reduce [7] to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes. However, their approach still relies on an event log that is initially stored at a central location. Subsequent work [12] extends their approach to the Flexible Heuristics Miner [23] and evaluates the performance of the Map-Reduce based Alpha Miner and the Flexible Heuristics Miner. While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research places dual emphasis on scalability, i.e, processing data directly at the source, and privacy preservation by only storing data aggregates at a central location.
Gatta et al. [14] address privacy concerns in the healthcare domain: Here, each hospital center aggregates their own event data into an FM, such that the others only see the aggregated versions of the data. Just as other works [2, 11, 12], they merge the FMs subsequently. Overall, all these collect event logs at one or a few locations. In EdgeAlpha, in contrast, each node collects a local event log and only shares aggregates, i.e., partial FMs, with the central location. As a result, EdgeAlpha increases scalability and reduces privacy risks.
4 Design
@@
4.1 Assumptions on Events and Event Logs
@@
In adherence to common practice, we assume events with the attributes case identifier, activity name and timestamp, whereas the case identifier is unique for each case. Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log. Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log. As we operate on distributed event logs, we – without loss of generality – assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
4.2 EdgeAlpha Algorithm
First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing EdgeAlpha algorithm. Without loss of generality, and to ease simplicity and comprehensibility, we assume each node to detect exactly one type of activity. Note, that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types. Furthermore, we assume that all nodes can communicate reliably and without significant delays, i.e., use common communication technologies and protocols such as cellular connectivity, Wi-Fi or Bluetooth.
@@ 4.2 EdgeAlpha Algorithm
EdgeAlpha as algorithm has two phases: Phase 1, initiated upon event triggering, and Phase 2, activated upon a process model request. In phase 1, whenever a node records an event, it finds the predecessor of the event by communicating to other nodes in the system and, once it determines the predecessor, it notes the directly follows relation in its local FM. As a result, every node continuously updates its local, partial FM over time. Phase 2 is activated whenever an entity requests a process model. Then all nodes send their partial FM to a central location which then constructs a FM from these partial ones. Overall, phase 2 is activated on demand and triggered by outside events, while phase 1 operates continuously. In the following, we define both phases. For simplicity, we begin by introducing the basic version of Phase 1, which, for each event, sends a message to each node in the system. Later, we introduce a more efficient variant and event batching.
4.2.1 Phase 1 – Event Ordering and Partial Footprint Matrix Construction.
@@
Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing. In this specific setting, event 1 is a start event. Thus, after the node senses the event, it queries for the predecessor event, to which it does not get any responses. Thus, it sets a flag that it detected a start event. For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response including event 1. It increments the value in the corresponding cell of its local FM.
@@
When an event $e$ on node $i$ is triggered, node $i$ executes the following steps, see Figure 2 for an example. First, node $i$ stores event $e$ locally, including its case identifier, the activity name and timestamp. Then, it sends all other nodes a message containing $e$ to determine the predecessor event. When receiving the request from $i$, a node checks its storage for events in the same case that have no successor yet and whose timestamps are smaller than the timestamp of $e$. If such an event exists, the node responds with the potential predecessor event data. If there are multiple responses, node $i$ picks the event with the latest timestamp, that is still smaller than the own one, as predecessor. We denote the predecessor node as node $j$. Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$. Subsequently, node $j$ saves $e$ as the successor event of its own event.
The case where there are multiple events to choose from, indicates communication delays. If delays result in incorrectly set predecessors or successors, the nodes rectify the stored events once the correct event is identified and update the counts in the successor’s partial FM accordingly. This is why nodes store a partial FM that not only encodes whether a direct succession has occurred, but also count the occurrences.
@@
Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessors in its FM. Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities. The value at index $\ell\in\{0,\dots,n-1\}$ describes how often $k$ directly follows $\ell$ in all of the cases, formally $\ell>_{L}k$. It is necessary to increment $FM^{k}_{\ell}$ by $1$ each time $k$ detects $\ell>_{L}k$ and subtract $1$ if a direct succession gets fixed. If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event. A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it directly.
4.2.2 Phase 2 – Requesting Process Data.

### v3

# EdgeAlpha: Distributed Process Discovery at

@@ Abstract
Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT) in a smart home, factory or hospital setting. The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs. Further, privacy considerations often prevent data collection at a central location in the first place. To address this challenge, this paper introduces EdgeAlpha, an algorithm for distributed process discovery operating directly on sensor nodes on a stream of real-time event data. Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded. From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model. Our analysis shows that even with over 600 activities, like in the Hospital Log, the average number of predecessors per activity remains below 7. This enables efficient determination of predecessors for each event, reducing the communication overhead by up to 96% compared to querying all nodes and enhancing scalability. Additionally, our analysis demonstrates that the number of queried nodes stabilizes after relatively few events, and batching predecessor queries in groups of 40 reduces the average queried nodes per event to less than 2.5%. EdgeAlpha ensures that the resulting process model is identical to the original Alpha Miner.
Index Terms:
@@
Distributed Process Mining, Event Data Stream, In-Network Processing, Scalability
I Introduction
@@
In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from Internet of Things (IoT) sensors, for example, in a smart manufacturing, smart home or smart health setting [1, 2, 3, 4]. However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [5, 6, 7]. This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [8]. Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place. For example, patient data in hospitals, personal data of workers, data in smart home setups, or company secrets in cross-organizational processes are commonly not shared for analysis [9, 10, 11]. We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
@@
Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes them in a distributed fashion directly at the source (right), and only exchanges data aggregates, increasing scalability and privacy.
In this paper, we introduce EdgeAlpha, an algorithm for distributed process mining based on the Alpha Miner [12]. Utilizing the constrained compute and storage resources of sensor nodes, further referred to as nodes, we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location. In contrast to Gatta et al. [9] which enable cross-organizational process mining through the transmission of data aggregates, EdgeAlpha distributes the computation of the footprint matrices (FMs) to the sensor-node level. This approach allows the collection of aggregated data within an organization that would otherwise not be obtainable in raw form due to, for example, too high data velocity or data protection regulations.
@@
For each event that a node in EdgeAlpha detects, it locally stores the event as well as the predecessor and successor of this event and updates a local FM accordingly, see Fig. 1. Thus, over time, each node constructs a partial FM that represents the view of this node on the process. To derive the complete process model, one of possibly multiple (central) entities queries the nodes for their partial FMs and combines these into a complete matrix. EdgeAlpha enables (a) scalable mining, as a node only interacts with its predecessors after each event and, when queried, only exchanges aggregates, i.e., partial FMs, with a central entity and (b) privacy-preserving process mining, as nodes only store their own as well as predecessor and successor events.
@@
To determine an event’s predecessor event, the node that records this particular event potentially has to query all other nodes, which can lead to large communication overheads. In this paper, we show that activities of real-world processes statistically only have a few predecessors activities, which further only rarely change throughout the lifetime of a process. Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [13] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. As a result, for each event, a node can very efficiently determine the predecessor reducing the communication overhead by up to 96% compared to querying all nodes in the network and thereby increasing scalability. Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events. Further, by batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes. In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor and a baseline that immediately queries the correct node. Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner and show that in four out of five datasets an intermediate result FM already complies with the FM of the Alpha Miner by more than 90% after less than 5% of the events of the datasets.
EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [14].
@@
  • We demonstrate that EdgeAlpha’s intermediate results quickly converge to the final result of the Alpha Miner.
  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach and by 97.5% when batching queries with a batch size of 40.
  • We ensure that the nodes only share aggregated event data with a central entity.
  • We implement EdgeAlpha and publish its code open-source [^1] Available at: redacted..
The remainder of the paper is structured as follows. Section II introduces the necessary background. Next, Section III presents related work on distributed process discovery. Afterwards, we introduce the design of EdgeAlpha in Section IV. In Section V, we evaluate EdgeAlpha. We conclude the paper in Section VI.
II Preliminaries and Background
In this section, we provide the required preliminaries and background. We provide definitions of direct successions and footprint matrices and recapitulate the steps of the Alpha Miner Algorithm [12] that are relevant to our approach.
@@ II Preliminaries and Background
Activity $b$ directly succeeds $a$ in event log $L$, i.e. $a>_{L}b$, if and only if there exists a trace $\sigma=(t_{0},\dots,t_{n-1})\in L$ and $j\in\{0,\dots,n-2\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
@@
Next, we recapitulate the initial steps of the Alpha Miner before later in the paper introducing a distributed variant. First, the algorithm identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$. Then, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and their end activities, respectively. Subsequently, the algorithm determines all direct successions in $L$ which can be stored in an FM. This FM is then the basis for the further calculations in the algorithm and the resulting Petri net.
A FM stores information regarding the direct succession relationships between activities within the event log. For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is an element of $\{0,\dots,|T_{L}|-1\}$. This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
@@
For our design, we define partial footprint matrices containing only one specific column: For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
III Related Work
Van der Aalst [18] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log. Here, models rather than FMs are created in a distributed manner. Evermann and Assadipour [19] leverage Map-Reduce [20] to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes. However, their approach still relies on an event log that is initially stored at a central location. Subsequent work [21] extends their approach to the Flexible Heuristics Miner [22] and evaluates the performance of the Map-Reduce-based Alpha Miner and the Flexible Heuristics Miner. While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research emphasizes scalability, i.e., processing data directly at the source and privacy preservation by only storing data aggregates at a central location.
Gatta et al. [9] address privacy concerns in the healthcare domain: In their work, each hospital center aggregates its own event data into an FM, such that the others only see the aggregated versions of the data. Like other works [18, 19, 21], they subsequently merge the FMs. In contrast, EdgeAlpha distributes the FM computations to the sensor nodes that record the events, enabling data aggregation at an earlier stage. This allows organizations to collect and utilize data in aggregated form that would otherwise be restricted in its raw form due to data protection regulations. As a result of processing events directly at their sources, EdgeAlpha increases scalability and reduces privacy risks.
IV Design
@@
IV-A Assumptions on Events and Event Logs
@@
In adherence to common practice, we assume events with the attributes case ID, activity name, and timestamp, whereas the case ID is unique for each case. Our evaluation shows these assumptions are both realistic and applicable in numerous real-world scenarios. For example, in a hospital each patient is commonly assigned a fixed ID that can serve as a case ID and the events such as blood tests or ultrasound are predefined activities. Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log. Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log. As we operate on distributed event logs, we assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
IV-B EdgeAlpha Setting
First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing the EdgeAlpha algorithm. We employ EdgeAlpha in typical IoT application settings: A network of sensor nodes that are, for example, integrated into machines in a smart factory or smart health setting. These sensors can communicate with each other, either directly or via the Internet. Further, we assume that a membership management allows nodes to gain knowledge of other nodes and how to reach them, as common in IoT applications. Together, these allow a node to find the processor and successor of an event. Finally, we assume that their clocks are well synchronized, using, for example, established clock-synchronization algorithms such as NTP [23]. This is essential, as accurate timestamps ensure proper event ordering. Without loss of generality, and to ease simplicity and comprehensibility, we assume each node detects exactly one type of activity. Note that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types.
IV-C EdgeAlpha Algorithm: Overview
@@
EdgeAlpha consists of two phases: Phase 1 is initiated upon event triggering, and Phase 2 is activated upon a process model request. In Phase 1, nodes communicate to determine predecessors when recording an event and note their directly-follows relations in their local FMs. As a result, every node continuously updates its local, partial FM. Other than Phase 1, Phase 2 is activated on demand whenever an entity requests a process model. Then, all nodes send their partial FMs to a central location, which constructs an FM. We define both phases, starting with the basic version of Phase 1, which sends a message to all nodes in the system for each event, followed by a more efficient variant and event batching.
IV-D Phase 1 – Event Ordering & Partial Footprint Matrices
@@
Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing. In this specific setting, event 1 is a start event. Therefore, after the node detects the event, it queries for the predecessor event, for which it does not get any responses. Thus, the node sets a flag that it detected a start event. For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response, including event 1. It increments the value in the corresponding cell of its local FM.
@@
When an event $e$ on node $i$ is triggered, node $i$ executes the following steps; see Fig. 2 for an example. First, node $i$ stores event $e$ locally, including its case ID, the activity name and timestamp. Then, it sends all other nodes a message containing $e$ to determine the predecessor event. When receiving the request from $i$, a node checks its storage for events in the same case that have no successor yet and whose timestamps are smaller than the timestamp of $e$. If such an event exists, the node responds with the potential predecessor event data. If there are multiple responses, node $i$ picks the event with the latest timestamp that is still smaller than the own one, as predecessor. We denote the node with the predecessor event as node $j$. Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$. Subsequently, node $j$ saves $e$ as the successor event of its own event.
If a node has to choose its predecessor from several events, it indicates communication delays which may cause incorrectly set predecessors. However, this is automatically rectified during the run of the algorithm as soon as the node that belongs to the incorrectly set predecessor event receives a request from the node of its correct successor. The node sets the correct successor and informs the node corresponding to the incorrect successor. This one starts the predecessor search again. The node rectifies the stored events once the correct event is identified and updates the counts in its partial FM accordingly. For this reason nodes store a partial FM that not only encodes whether a direct succession has occurred but also counts the occurrences. The rectification process works due to a single start event and because the nodes compare the timestamps of set successor events with the timestamps of incoming predecessor-search-requests. As soon as a node has more than one successor in a single trace the incorrectly set one is informed and an total order of events is established.
@@
Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessor activities in its FM. Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities. The value at index $\ell\in\{0,\dots,n-1\}$ describes how often $k$ directly follows $\ell$ in all of the cases, formally $\ell>_{L}k$. It is necessary to increment $FM^{k}_{\ell}$ by $1$ each time $k$ detects $\ell>_{L}k$ and subtract $1$ if a direct succession gets fixed. If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event. A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it immediately.
EdgeAlpha’s design further incorporates a sliding-window-principle for discarding old data in order to respond to the requirements of the resource-restricted sensor nodes. With this principle, as soon as a certain memory size has been reached or a certain time has been exceeded, the node discards the oldest data when adding new data. However, the sliding window approach only gets beneficial with a vast amount of data. The event logs we consider in the evaluation do not fill enough memory to utilize the sliding window.
IV-E Phase 2 – Requesting Process Data